OpenStudy (anonymous):

True or False? Y varies directly with x and y = -9 when x = 3. When y = 27, the value of x is 9.

OpenStudy (callisto):

Y varies directly with x => y = kx y = -9 when x = 3 => -9 = k(3) Solve k Then, solve x when y = 27

OpenStudy (anonymous):

I dont know

OpenStudy (callisto):

Which part?

OpenStudy (anonymous):

y= 27 the whole formula

OpenStudy (callisto):

Y varies directly with x => y = kx ^ Do you understand this part?

OpenStudy (anonymous):

WAIT, -9 = k(3)

OpenStudy (anonymous):

Then do I times 3 on both sides

OpenStudy (callisto):

Nope, to solve k, you need to divide 3 on both sides.

OpenStudy (callisto):

Oh wait, another little trick

OpenStudy (callisto):

-9 = k(3) ^ Multiply -3 to both sides, so that you can get 27 on the left. ( y=27 as given in the question) Then, see if the coefficient of k is 9. If it is, then it is true, and vice versa.

OpenStudy (anonymous):

OH ok thanks

OpenStudy (callisto):

Welcome :)